# When recalling in-context, Transformers are not SSMs

### v1

@@ Acknowledgements
We would like to thank Alberto Bietti for his helpful feedback. The authors thank Schmidt Sciences for the support received through the AI2050 program. Antonio Orvieto also acknowledges the financial support of the Hector Foundation.
@@ A.2 Experimental details
In this section, we describe the experimental setup used throughout our study. Clearly outlining these details is crucial for interpreting the results presented in subsequent sections. Our implementation is inspired by methodologies from Zoology  [Arora et al., 2023].
@@
Each model is tested across six model dimensions: $64$, $128$, $256$, $512$, $1024$ and $2048$. Additionally, models are implemented in two configurations: 1-layer and 2-layer. Notably, a “layer” in our context refers to the concatenation of two blocks: a sequence mixer (e.g., attention, RWKV, etc.) followed by an MLP. Thus, a 1-layer model consists of two blocks, aligning with the terminology used in prior work  [Arora et al., 2023, Olsson et al., 2022]. Positional information is used only in attention and Based.
@@
A.3 Full Tables
Figure 10: Performance of 2-layers models. We report the official results [^6] Mamba was not included in the official work but some experiments, with different settings compared to ours, are documented in the blog post. ( green stars) and the replication running the original code of [Arora et al., 2023] ( dotted blu line). While for replication, we used the learning rates grid by Arora et al. [2023], we note here that, due to high sensitivity to the learning rate (Fig, 1), tuning drastically affects performance. In solid orange, we provide results with a finer grid (cf. Fig.1). Careful tuning of the learning rate gives a general improvement in the performance of recurrent models. This becomes especially crucial in Mamba, where the task becomes solvable at high sequence lengths >>>> hidden size. The results show the mean and relative max-min errors for 3 seeds. Attention always solves the task (all curves overlap).

### v2

@@ Acknowledgements
We would like to thank Alberto Bietti for his helpful feedback. The authors thank Schmidt Sciences for the support received through the AI2050 program. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Destiny Okpekpe. Antonio Orvieto also acknowledges the financial support of the Hector Foundation.
Ethics Statement
As our research focuses on a foundational analysis of sequence models using synthetic data we do not foresee any direct ethical concerns arising from our methods or findings.
Reproducibility Statement
To ensure the full reproducibility of our experimental findings, all of our code is made publicly available. Our implementation builds upon the open-source codebases provided by  (Arora et al., 2023) and  (Jelassi et al., 2024). All specific hyperparameters, architectural details, and training configurations for each model and task are comprehensively documented in Appendix A.2. We believe these measures provide everything necessary for the direct replication of our results.
Acknowledgment of AI-assisted Tools
AI-assisted editing tools were used to check grammar.
@@ A.2 Experimental details
In this section, we describe the experimental setup used throughout our study. Clearly outlining these details is crucial for interpreting the results presented in subsequent sections. Our implementation is inspired by methodologies from Zoology  (Arora et al., 2023).
@@
Each model is tested across six model dimensions: $64$, $128$, $256$, $512$, $1024$ and $2048$. Additionally, models are implemented in two configurations: 1-layer and 2-layer. Notably, a “layer” in our context refers to the concatenation of two blocks: a sequence mixer (e.g., attention, RWKV, etc.) followed by an MLP. Thus, a 1-layer model consists of two blocks, aligning with the terminology used in prior work  (Arora et al., 2023; Olsson et al., 2022). Positional information is used only in Attention and Based.
@@
A.3 MQAR performance of 2-layer models
Figure 11: Performance of 2-layers models. We report the official results [^7] Mamba was not included in the official work but some experiments, with different settings compared to ours, are documented in the blog post. ( green stars) and the replication running the original code of (Arora et al., 2023) ( dotted blu line). While for replication, we used the learning rates grid by Arora et al. (2023), we note here that, due to high sensitivity to the learning rate (Fig, 1), tuning drastically affects performance. In solid orange, we provide results with a finer grid (cf. Fig.1). Careful tuning of the learning rate gives a general improvement in the performance of recurrent models. This becomes especially crucial in Mamba, where the task becomes solvable at high sequence lengths >>>> hidden size. The results show the mean and relative max-min errors for 3 seeds. Attention always solves the task (all curves overlap).
A.4 MQAR performance of 1-layer models
Figure 12: Performance of 1-layer models on MQAR. We show how for recurrent models, scaling the width boosts performances. On the contrary, Attention models cannot solve the task anymore as in the 2-layer setting, and performances are unaffected by the scaling in width. The results show the mean and relative max-min errors after 3 runs with different seeds.
A.5 MQAR performance of ablated 1-layer models
Table 3: Performance of ablated 1-layer architectures on MQAR. Convolutions allows a 1-layer Transformer to reach perfect accuracy, while Mamba preserve its performance in all ablations.
Model Solves MQAR
Attention $2\%$
Attention + Conv on QKV $99\%$
Attention + Conv on K $99\%$
Attention + Conv on V $99\%$
Attention + Conv on Q $2\%$
Mamba $99\%$
[…]
S6 + MLP (Mamba as a Transformer) $98\%$
Table 4: Effect of various PEs in MQAR. Performance are comparable and aligns with the fact that recurrent model already encode positional information in their recurrence
Architecture No PE Absolute PE Relative PE Learned PE
Mamba $9\pm 0.01 0.99±0$$9\pm 0.01$ $9\pm 0.01 0.99±0$$9\pm 0.01$
Hyena $0\pm 0.11 0.33±0$$3\pm 0.11$ $0.31\pm 0.04$ $0.29\pm 0.08$
RWKV $0.19\pm 0.09$ $2\pm 0.09 0.22±0$$2\pm 0.09$ $0.24\pm 0.13$
H3 $0.24\pm 0.15$ $0.25\pm 0.10$ $0.27\pm 0.07$ $0.28\pm 0.11$